OpenStudy (anonymous):

This is a problem that i want to solve but i can't really answer it. It's a sample problem from my book: The density of a copper is 8.89 is 8.89 g/cm^3 and the ρ=1924μΩ-cm. Determine the length(m) and resistance of 12kg of copper; the cross-section of which in 2.09mm^2.

OpenStudy (ash2326):

Copper wire can be considered as a cylinder. Here we have cross section given as 2.09 mm^2 To find length, we need to find the volume We have the density as 8.89 g/ cm^2 Mass= 12 kg Could you find the volume? Do remember to change mass in kg to grams, this will give you volume in cm^3

OpenStudy (anonymous):

so 12000 g?

OpenStudy (ash2326):

Yeah, now find volume Volume= mass/density

OpenStudy (anonymous):

density hai mass hai volume= length into cross section you can find length

OpenStudy (anonymous):

1,349.83 cm^2 ?

OpenStudy (anonymous):

@ash2326 @theyatin what do you think?

OpenStudy (ash2326):

Yeah you're right. Sorry My battery was out of juice

OpenStudy (anonymous):

than after i got the volume, whats next? getting the length?

OpenStudy (anonymous):

*then

OpenStudy (ash2326):

\[\text{Volume}= \text{Area of Cross section} \times \text{length}\]

OpenStudy (anonymous):

length=volume/ area of cross section?

OpenStudy (ash2326):

Yeah, remember to convert cross section into cm^2, it's given in mm^2 Then you'll get length in cm

OpenStudy (anonymous):

length= 1,349.83 cm^2/0.209cm^2 ??

OpenStudy (ash2326):

1 mm = 1/ 10 cm \[1 mm^2= \frac{1}{10\times 10} cm^2\]

OpenStudy (anonymous):

.209 cm^2 ?

OpenStudy (ash2326):

area of cross section=2.09 mm^2=2.09/100 cm^2

OpenStudy (anonymous):

so 0.0209 cm^2 ?

OpenStudy (ash2326):

yeah now you find length

OpenStudy (anonymous):

length= 1,349.83 cm^2/0.0209cm^2 ?

OpenStudy (anonymous):

64,585.17 ?

OpenStudy (anonymous):

what would be the unit?

OpenStudy (ash2326):

\[\frac{cm^3}{cm^2}\]???

OpenStudy (anonymous):

oh i get it. its cm. cause i saw both cm^2

OpenStudy (ash2326):

good, now you have the length. You need to find the resistance. Do you have any idea, how to find that?

OpenStudy (anonymous):

resistance? R=v/i ?

OpenStudy (ash2326):

that's right but there is one more relation relating resistance with the property of material and the dimensions of the wire. \[R=\rho(\text{micro ohms-cm}) \times \frac{\text{Length (in cm)}}{\text{Area of cross section(in cm^2)}}\] From this you'll get resistance in micro ohms

OpenStudy (anonymous):

R=ρ(micro ohms-cm)×3090199.522 cm ?

OpenStudy (anonymous):

then multiply it to 1924 ?

OpenStudy (ash2326):

Yeah:D

OpenStudy (anonymous):

5945643879 ?

OpenStudy (ash2326):

Yeah this is right but this in micro ohms. Divide by 10^6 to get resistance in ohms

OpenStudy (anonymous):

why divide it by 10^6 ?

OpenStudy (ash2326):

Because resistance we found is in micro ohms \[1 micro ohm=\frac{1}{10^6} \text{ohms}\]

OpenStudy (anonymous):

oh okay thanks so the answer will be 5945.54 ohms ?

OpenStudy (ash2326):

Yeah:D

OpenStudy (anonymous):

then whats next? ;D

OpenStudy (ash2326):

Yeah that's all we found length and resistance Wait a min, we need to find length in meters. For that divide length in cm by 100 as 1 meter= 100 cms

OpenStudy (anonymous):

so that must be 645.8517 ?

OpenStudy (ash2326):

64585.17/100=>64.581517 meters

OpenStudy (anonymous):

no its 645.8517

OpenStudy (ash2326):

yeah, you're right. Sorry :(

OpenStudy (anonymous):

its alright :) then whats next? we will repeat it again? because we got the wrong length?

OpenStudy (ash2326):

nope, we found the resistance correctly. question asked length in meters that's why we changed it

OpenStudy (anonymous):

oh thanks a lot. :D a massive thank you! God bless :D

OpenStudy (ash2326):

welcome:D